what is the length and width of a triangular sign if the length exceeds the width by 15 inches and the perimeter is 158 inches?
Answer by htmentor(1343)

Since the perimeter is the sum of all the sides, 2l + w = 158 (1)
Also l - w = 15 (2)
Solving for l in (2) and substituting in (1) we have 2*(w+15) + w = 158
This reduces to 3w + 30 = 158, or w = 128/3 = 42 2/3
Putting this into (2) we have l - 42 2/3 = 15 or l = 57 2/3.